Case Study: In addition to regular gyms, nontraditional workout concepts and centers such as Kosama are increasing in popularity. Kosama is a franchise opportunity that offers members the opportunity to improve their health and fitness level. To learn more about the company visit kosama.com.

Part 1, Section 1: Assume the following revenue and cost break-down.

Revenue:  

-Monthly membership fee = $28.

Costs:

-General fixed operating expenses = $3,975 per month.

-Equipment Lease = $410 per month.

-Towel service = $.50 per member based on volume.

-Mixed costs are equal to $275 per/month (fixed) plus $1.10 per membership sale (variable).

-Total variable costs are not known.  

-Estimated number of members required to break even is between 255 and 275 members per month.  

Part 1, Section 2: Using the information from section 1. What would monthly sales in members and dollars have to be to achieve a target net income of $14,500 for the month? What is the margin of safety in dollars? Show your work and all calculations.

Part 1, Section 3: Discuss how cost structure, relevant range, margin of safety, cost behaviors, and CVP apply to an investment in the franchise. How do you plan to use this in order to manage the business and plan for profitability? What type of internal accounting reports would you prepare? Why?

Part 1, Section 4: Assume you decide to invest in the franchise. Provide a description and estimates in dollars for sales, variable and fixed expenses. Explain how you determined each number and provide a written list of assumptions.
Answer #1

Firstly, the above scenario is a service based revenue that has been incurred. Calculation of monthly sales:

Cost of services-net service revenue, which is: $4,661.60-$14,500 = $9,838.40 is the gross profits. The, gross profits+operating expenses, which is = $9,838.40+$3,975 = $13,813.40. Therefore, operating income would be, gross profits-expenses, which is $9,838.40-13,813.40 = 3,975 is the operating income.

Section 2:Monthly membership fee is $28 & net income monthly target is $14,500. $14,500/$28 = $517.85 would be monthly sales. Margin of safety is break even sales-projected sales , which is; $28 *530 = $14,840( break even sales) - $14,500 = $340.

Section 3: Profitability is calculated as follows: revenue-cost of goods sold/revenue*100%

= $14,500-$4,662.70/$14,500*100% = $0.678 % would be the profit margin. Some of the internal accounting reports would be: cash reports, status reports, financial reports, budget books & board reports. These reports are prepared to be aware of the company's activities & to check on the financial health of the company & to make some key decisions.